OpenStudy (anonymous):

Solve each equation. Check for extraneous solutions. |3x+5|=5x+2

OpenStudy (anonymous):

Use the definition of absolute value. Either (3x+5)=5x+2 or -(3x+5)=5x+2 Solving the first equation gives 2x=3 or x=3/2 Solving the second equation gives 8x=-7 or x=-7/8. Checking both answers in the original equations, it looks to me like -7/8 is extraneous, since the left side would be + and the right side would be negative. So x=3/2.

OpenStudy (anonymous):

Verifying it graphically confirmed that there is only one solution, by the way.